OpenStudy (anonymous):

9 divided by 28 1/4?

OpenStudy (phi):

can you change 28 1/4 to an improper fraction ?

OpenStudy (anonymous):

Yes.

OpenStudy (phi):

the first step is change 28 1/4 to an improper fraction next step: instead of dividing "flip" the fraction and multiply example: 2 divided by 1/3 is the same as 2 times 3 (3 is 1/3 flipped)

OpenStudy (anonymous):

So, 28 times 4?

OpenStudy (phi):

yes the top becomes 28*4 + 1 the bottom stays 4

OpenStudy (anonymous):

So 9 divided by 112

OpenStudy (phi):

in case you did not know \( 28 \frac{1}{4} \) is the same as 28 + \(\frac{1}{4} \) or \[ \frac{28\cdot 4}{4} + \frac{1}{4} = \frac{4 \cdot 28+1}{4} \]

OpenStudy (anonymous):

Okay thanks

OpenStudy (phi):

so you have 9 divided by 113/4 flip the fraction and multiply 9 * 4/113

OpenStudy (anonymous):

I came up with 38/113 ?

OpenStudy (phi):

is your calculator broken ?

OpenStudy (anonymous):

no

OpenStudy (phi):

how did you get 38 ?

OpenStudy (anonymous):

I have no idea ..

OpenStudy (phi):

your problem is \[ \frac{9}{1} \cdot \frac{4}{113} \] to multiply fractions, you multiply top times top and bottom times bottom

OpenStudy (anonymous):

I would 36 and 113 ?

OpenStudy (anonymous):

36/113

OpenStudy (phi):

yes, 36/113 looks good

OpenStudy (anonymous):

Cool thanks